OpenStudy (anonymous):

Find the product. 8y^3(-3y^2) is it .. -24y^5?

mathslover (mathslover):

yes u r right again

mathslover (mathslover):

\[\large{8y^3\times -3y^2}\] \[\large{8\times -3 \times y^{3+2}}\] \[\large{-24 y^5}\]